OpenStudy (anonymous):

study help::: * dont need answer need to learn how to solve for this type of problems... Six differently colored balls (red, blue, green, orange, purple, and white) are placed in a basket. Without looking, three balls are removed. What is the total number of combinations that include a red ball?

OpenStudy (anonymous):

I'll help in one sec.

OpenStudy (anonymous):

Ok so basically your determining probabilities, right?

OpenStudy (anonymous):

yes

OpenStudy (kinggeorge):

If we ignored the fact that one of the balls has to be red, how many combinations would there be of three balls?

OpenStudy (anonymous):

two

OpenStudy (kinggeorge):

There would actually be quite a few more. We could have red, green and blue. Or maybe red, blue, and orange. Perhaps I have red, blue, and white. To find the total number of combinations, let's start with one. How many combinations of a single ball are there?

OpenStudy (anonymous):

six

OpenStudy (kinggeorge):

Exactly. Now, how many combinations of 2 balls? Remember that there should be more than three combinations.

OpenStudy (anonymous):

hmm 9?

OpenStudy (kinggeorge):

Look at it this way. You have 6 options for the first ball you can choose. Now, since you've already chosen one, you only have 5 balls left to choose from. So you should have \[6\cdot5=30\]Ways to choose to balls. Does this make sense?

OpenStudy (anonymous):

yeah im getting multiply by watever u have left if there were more numbers id multiply them all

OpenStudy (kinggeorge):

There's one more thing I forgot to mention though. The way we calculated this, if we chose a red ball and then the blue ball, we would have a different set than if we chose the blue ball then the red ball. So really, we want \(6\cdot5\) divided by the number of ways we can order the two balls. since there's only two balls, they can only be ordered in two ways. Thus, we actually want \[{6\cdot5 \over 2}=15\] Sorry about the confusion. By the way, how much have you talked about combinations in your class so far?

OpenStudy (anonymous):

i study online and theres only quizzes about this type of problems got a tst tomorrow and im sure this type of problems will be there and i got you on watch u just said about dividing

OpenStudy (anonymous):

what*

OpenStudy (kinggeorge):

Basically, what I did there by dividing by 2, is using the formula\[{n! \over k!(n-k)!}\]With \(n=6\) and \(k=2\). This formula gives you the number of combinations of \(k\) things from an overall group of \(n\) things. Have you seen this formula before?

OpenStudy (anonymous):

yes

OpenStudy (kinggeorge):

This is what you want to use for combinations. Using this formula, can you solve for the case of choosing 3 balls out of 6 total?

OpenStudy (anonymous):

n=6 k=3? i mean using these?

OpenStudy (kinggeorge):

Exactly.

OpenStudy (anonymous):

6/9

OpenStudy (kinggeorge):

Remember that \[6! =6\cdot5\cdot4\cdot3\cdot2\cdot1=720\]And that \[3! \cdot(6-3)!=3!\cdot3!=3\cdot2\cdot1 \cdot 3\cdot2\cdot1=36\]

OpenStudy (anonymous):

ok il write that down

OpenStudy (kinggeorge):

So in the end, you should get \(720/36 =20\) combinations with 3 balls from 6 original balls.

OpenStudy (kinggeorge):

Now we have to get back to the original problem of choosing three balls from six if one of those three must be red. Let's say you've already chosen the red ball. That leaves us with 5 balls left to choose from, and only two spots left. Thus, we're asking how many combinations of 2 balls can be chosen from a set of 5 original balls. This problem is solved by the formula I wrote above.

OpenStudy (anonymous):

thank u kinggeorge for bearing with me